# Task-Agnostic Federated Continual Learning via Replay-Free Gradient Projection

### v1

@@ 1 Introduction
Federated learning (FL) (McMahan et al., 2017), where client devices collaboratively train a global model without sharing private data, has emerged as a compelling alternative to centralized learning. Most FL systems assume static local datasets and a single inference task per client. In practice, however, devices (e.g., phones, smart glasses) often collect data for multiple evolving tasks and must fine-tune the model over time. Limited storage further forces clients to discard old data, creating a continual learning (CL) scenario where models must adapt to new tasks without having access to prior ones. This setting exacerbates catastrophic forgetting (McCloskey & Cohen, 1989), i.e., deterioration of the model’s performance on previously learned tasks. In federated systems that operate under continual learning, this problem is further intensified by the data heterogeneity across client devices.
Many studies have explored adapting conventional continual learning schemes to federated settings, typically falling into three categories: (1) replay-based methods (Dong et al., 2022; Liu et al., 2023; Dai et al., 2023; Li et al., 2024c; a); (2) generation-based methods (Qi et al., 2023; Zhang et al., 2023; Tran et al., 2024; Liang et al., 2024; Yu et al., 2024); and (3) regularization-based methods (Yoon et al., 2021; Ma et al., 2022; Li et al., 2024b; Lee et al., 2024). Each faces challenges in FL: replay methods store old examples, risking privacy or exceeding storage limits; generation methods require server-side generative models, slowing aggregation; and regularization methods often impose significant local compute overhead. Recently, gradient projection methods such as GPM (Saha et al., 2021) have proven effective in centralized continual learning by projecting new task gradients onto subspaces orthogonal to those of previously learned tasks. However, GPM assumes centralized access to activation statistics, which renders it impractical in federated settings. While FOT (Bakman et al., 2024) adapts GPM to FL by collecting high-dimensional embeddings from clients, this approach incurs significant communication overhead and raises potential privacy concerns.
@@ 2.1 Federated Continual Learning
Federated continual learning (FCL) tackles the challenge of continuously learning from decentralized data while preserving performance across sequential tasks. An early FCL approach, FedWeIT (Yoon et al., 2021), decomposes model parameters into task-generic and task-specific components, focusing on a task-incremental setting where the task ID is known during inference. CFeD (Ma et al., 2022) relies on knowledge distillation using a surrogate dataset shared between the server and clients. GLFC (Dong et al., 2022; 2023) addresses catastrophic forgetting by combining class-aware gradient compensation with class-semantic relation distillation, but relies on storing examples from previous tasks. Subsequent works (Liu et al., 2023; Dai et al., 2023; Li et al., 2024c; a) reduce the size of the replay cache, yet remain reliant on stored samples.
@@
Recently, several FCL methods have leveraged generative models to replace real examples in memory with synthetic data. FedCIL (Qi et al., 2023) employs a GAN with an auxiliary classifier to enable generative replay, thereby mitigating forgetting while aggregating global knowledge across clients. TARGET (Zhang et al., 2023) and MFCL (Babakniya et al., 2024) introduce data-free knowledge distillation that uses synthetic examples to transfer knowledge from a previously trained global model to client models. LANDER (Tran et al., 2024) builds on this idea by incorporating label text embeddings from pretrained language models as anchors, which enables the generation of more meaningful samples and enhances resistance to forgetting. While effective in mitigating forgetting, these approaches inherit significant drawbacks: training generative models is computationally expensive as image resolution increases and introduces new privacy risks (Liu et al., 2024).
@@
Overall, existing FCL methods face practical challenges in real-world FL deployments due to privacy concerns and resource constraints. Specifically, they often: (1) assume the task identity is known during inference; (2) store exemplars from previous tasks on the server; or (3) train a generative model to synthesize replay samples. In contrast, FedProTIP requires none of these assumptions. It is explicitly designed for task-agnostic inference, where task labels are unavailable, and achieves this without replay buffers, generative models, or auxiliary task classifiers. Instead, FedProTIP leverages lightweight subspace representations for both knowledge retention and task-identity prediction, enabling effective task-agnostic FCL. This focus connects to the broader literature on class-incremental learning (CIL). (Kim et al., 2022b) provides a theoretical perspective that decomposes the CIL problem into within-task classification and task-identity prediction, showing that both are necessary and sufficient conditions for strong performance. While centralized methods address the challenge of task-agnostic inference via out-of-distribution detection (Kim et al., 2022b; a), per-class classifiers or generative models (Zając et al., 2024; Van De Ven et al., 2021), or supervised contrastive learning with nearest-class-mean classifiers (Mai et al., 2021), none extend naturally to federated environments. Our work is the first to bring task-identity prediction into the FCL paradigm, providing a replay-free, privacy-preserving solution in the challenging task-agnostic CIL setting.
@@ 2.2 Gradient Projection in Continual Learning
Gradient projection methods (Zeng et al., 2019; Farajtabar et al., 2020; Chaudhry et al., 2020) for continual learning mitigate forgetting by updating model parameters in directions orthogonal to those associated with previous tasks, thereby eliminating the need to store raw data or train generative models. GPM (Saha et al., 2021) extends these approaches by extracting the bases of low-dimensional subspaces spanned by prior task representations and constraining new gradients to lie orthogonal to these subspaces. A series of follow-up works, including TRGP (Lin et al., 2022b), CUBER (Lin et al., 2022a), SGP (Saha & Roy, 2023) and DualGPM (Liang & Li, 2023a), relax the orthogonality constraints introduced in GPM to better balance stability and plasticity, which leads to improved continual learning performance. However, directly extending gradient projection methods to federated continual learning (FCL) is nontrivial due to the decentralized nature of data and limited communication budgets. FOT (Bakman et al., 2024) represents a recent attempt to adapt GPM to the FCL setting by requiring clients to share their raw feature embeddings with a central server to construct gradient subspaces. This approach introduces privacy risks and communication overhead, and its effectiveness is limited to scenarios where task identities are known at inference time – an impractical assumption in many real-world settings. In contrast, FedProTIP achieves state-of-the-art FCL performance using a communication- and computation-efficient projection scheme, without exposing raw embeddings or relying on task labels during inference.
@@ 3.2 Gradient Projection Memory
Gradient projection memory (Saha et al., 2021) is a replay-free CL scheme that requires storing only a set of core bases $\bm{\Phi}_{l}^{(1:t)}$ extracted from layer-wise activations after fine-tuning the model on $t$ tasks. Specifically, let $\mathbf{W}_{l}^{(t)}\subset\mathbf{W}^{(t)}$ denote the parameters of layer $l$ after training on task $t$, and let $\mathbf{a}_{l}^{(t)}\in\mathbb{R}^{d_{l}\times m}$ represent the input activations to layer $l$ for $m$ training samples $\mathbf{x}^{(t)}$, where $d_{l}$ is the dimensionality of the activations. By applying singular value decomposition (SVD), GPM extracts a set of orthonormal bases $\bm{\Phi}_{l}^{(t)}\in\mathbb{R}^{d_{l}\times r_{l}^{(t)}}$ that span the dominant subspace of task $t$ activations and aggregates them with the existing bases $\bm{\Phi}_{l}^{(1:t-1)}$. During training on the (t+1)(t{+}1)-th task, the parameter update for layer $l$, denoted $\Delta\mathbf{W}_{l}^{(t+1)}$, is projected onto the orthogonal complement of the subspace spanned by $\bm{\Phi}_{l}^{(1:t)}$,
@@ 4 Methodology
While GPM has proven effective in centralized continual learning, extending it to federated settings poses major challenges. FOT (Bakman et al., 2024) offers an early adaptation by having clients share layer-wise intermediate activations, which the server uses to extract core bases. However, this approach raises significant privacy concerns, as such activations can be exploited in gradient inversion attacks (Geiping et al., 2020; Chen & Vikalo, 2024). It also introduces substantial communication overhead due to the high dimensionality of the transmitted activations.
@@ 4.2 Extracting Local Core Bases
After completing projected gradient descent-based local training, clients follow the standard federated learning protocol by sending local model updates to the server for aggregation. The server then broadcasts the updated global model $\mathbf{W}^{(t)}$ to all clients, which proceed to perform local core basis extraction. Following the GPM strategy (Saha et al., 2021), each client $k$ samples $m$ examples from its local dataset $\mathcal{D}_{k}^{(t)}$, feeds them through the received model $\mathbf{W}^{(t)}$, and collects layer-wise intermediate activations. To reduce storage and communication overhead, we introduce a random activation sampling step: from the $m$ activations, a smaller subset of size $m^{s}\!\!\ll\!\!m$ is randomly selected, yielding $\mathbf{a}_{l}^{(t)}\in\mathbb{R}^{d_{l}\times m^{s}}$. These activations are projected onto the orthogonal complement of the previously learned feature subspace by subtracting their component along the existing bases,
@@ 4.4 Task Identity Inference via Subspace Relevance
In continual learning, the feature extractor (encoder) is fine-tuned across sequential tasks, while the decision layer $f_{1:t}(\cdot)$ expands as new tasks are introduced. For example, in class-incremental settings, the output dimensionality of the softmax layer grows with the number of classes. Prior works (Saha et al., 2021; Bakman et al., 2024) assume that the task identity $\tau$ is known at inference time, so predictions can be routed through the corresponding decision head $f_{\tau}(\cdot)$. However, this assumption is unrealistic in real-world deployments, where task labels are typically unavailable (Kim et al., 2022b).
@@ 5 Experiments
We evaluate FedProTIP on three standard continual learning benchmarks: CIFAR100 and ImageNet-R (Hendrycks et al., 2021) for class-incremental learning, and DomainNet (Peng et al., 2019) for domain-incremental learning. Comparisons are made against six baselines: FedAvg (McMahan et al., 2017), GLFC (Dong et al., 2022), LGA (Dong et al., 2023), TARGET (Zhang et al., 2023), FOT (Bakman et al., 2024), and LANDER (Tran et al., 2024). Following (Yurochkin et al., 2019), we simulate non-IID client distributions by sampling data partitions via a Dirichlet distribution with varying concentration parameter $\alpha$ (lower $\alpha$ implies greater heterogeneity). All methods use a ResNet-18 backbone pretrained on ImageNet-1K (He et al., 2016) and fine-tuned on each benchmark dataset. Additional results using ResNets and ViTs trained from scratch are provided in Appendix A.1.
Following prior work (Chaudhry et al., 2018), we evaluate performance using two standard metrics: average accuracy (ACC) and forgetting (FT), defined as
@@ 5.2 Robustness under Data Heterogeneity and Forgetting
As shown in Table 1, FedProTIP consistently outperforms all baselines across various values of the Dirichlet concentration parameter $\alpha$, which controls the degree of data heterogeneity (larger $\alpha$ corresponds to more IID-like partitions). As heterogeneity increases (i.e., smaller $\alpha$), client drift (Karimireddy et al., 2020) exacerbates catastrophic forgetting, degrading the performance of most methods. Despite this, FedProTIP remains robust. For example, on CIFAR100, while competing baselines like FOT and LANDER experience accuracy drops of 12% and 15%, respectively, when moving from IID to $\alpha=0.2$, FedProTIP sees only a 6% decline. Moreover, it sustains high accuracy while keeping forgetting near zero, demonstrating strong resilience to heterogeneous client updates. Beyond class-incremental learning, we evaluate robustness under overlapping task semantics using DomainNet in a domain-incremental setting, where all tasks share the same label space. In the task-agnostic scenario, both the baselines and FedProTIP (-t) use a single shared classifier. Under this setup, FedProTIP (-t) consistently outperforms all competing methods across varying levels of heterogeneity, demonstrating benefits that extend beyond task identity prediction by enabling more robust representation learning. In contrast, FedProTIP maintains separate classifiers for each task and uses task prediction to route test inputs. Although this approach yields slightly lower accuracy than FedProTIP (-t), which uses a shared classifier, FedProTIP remains competitive – it achieves the lowest forgetting across all settings and delivers accuracy second only to our FedProTIP (-t).
@@ 5.3 Ablation Studies
We investigate whether task-agnostic continual learning methods originally developed for centralized settings can be adapted to federated scenarios. Specifically, we consider the replay-free PEC (Zając et al., 2024), which assigns a separate classifier to each class, and the replay-based SCR (Mai et al., 2021), which uses a nearest-class-mean classifier. For SCR, we average class prototypes across clients at inference time while keeping replay data local, and use FedAvg for model aggregation. As shown in Table 5, both methods yield significantly lower accuracy under task-agnostic inference. While SCR outperforms PEC, it relies on clients retaining local data, which may violate privacy constraints. We also evaluate LODE (Liang & Li, 2023b), which decouples intra- and inter-task losses to implicitly support task-agnostic inference. Applied to the generative replay method TARGET, LODE offers only marginal gains ($2.50\%$), suggesting that naive loss decoupling is insufficient for robust performance in federated settings.
@@ A.1 Results on Different Models
We also evaluate our method on a different backbone, pre-trained ViT-B/16 (Steiner et al., 2021), as reported in Table 6. In this setting, CIFAR100 images ($3\times 32\times 32$) are resized to $224\times 224$ to match the ViT input resolution. We set the number of local epochs to $5$ and perform $20$ global rounds per task. For Target and Lander, we observed that generating synthetic images at the native $32\times 32$ resolution and subsequently applying resizing augmentation yields better performance, and we adopt this strategy in our experiments.

### v2

@@ 1 Introduction
Federated learning (FL) (fedavg), where client devices collaboratively train a global model without sharing private data, has emerged as a compelling alternative to centralized learning. Most FL systems assume static local datasets and a single inference task per client. In practice, however, devices (e.g., phones, smart glasses) often collect data for multiple evolving tasks and must fine-tune the model over time. Limited storage further forces clients to discard old data, creating a continual learning (CL) scenario where models must adapt to new tasks without having access to prior ones. This setting exacerbates catastrophic forgetting (mccloskey1989catastrophic), i.e., deterioration of the model’s performance on previously learned tasks. In federated systems that operate under continual learning, this problem is further intensified by the data heterogeneity across client devices.
Many studies have explored adapting conventional continual learning schemes to federated settings, typically falling into three categories: (1) replay-based methods (Dong2022GLFC; liu2023fedet; dai2023fedgp; li2024sr; li2024towards); (2) generation-based methods (qi2023FedCIL; zhang2023target; tran2024lander; liang2024dddr; yu2024overcoming); and (3) regularization-based methods (yoon2021fedweit; ma2022CFeD; li2024rehearsal; lee2024fedsol). Each faces challenges in FL: replay methods store old examples, risking privacy or exceeding storage limits; generation methods require server-side generative models, slowing aggregation; and regularization methods often impose significant local compute overhead. Recently, gradient projection methods such as GPM (saha2021GPM) have proven effective in centralized continual learning by projecting new task gradients onto subspaces orthogonal to those of previously learned tasks. However, GPM assumes centralized access to activation statistics, which renders it impractical in federated settings. While FOT (bakman2024FOT) adapts GPM to FL by collecting high-dimensional embeddings from clients, this approach incurs significant communication overhead and raises potential privacy concerns.
@@ 2.1 Federated Continual Learning
Federated continual learning (FCL) tackles the challenge of continuously learning from decentralized data while preserving performance across sequential tasks. An early FCL approach, FedWeIT (yoon2021fedweit), decomposes model parameters into task-generic and task-specific components, focusing on a task-incremental setting where the task ID is known during inference. CFeD (ma2022CFeD) relies on knowledge distillation using a surrogate dataset shared between the server and clients. GLFC (Dong2022GLFC; dong2023LGA) addresses catastrophic forgetting by combining class-aware gradient compensation with class-semantic relation distillation, but relies on storing examples from previous tasks. Subsequent works (liu2023fedet; dai2023fedgp; li2024sr; li2024towards) reduce the size of the replay cache, yet remain reliant on stored samples.
@@
Recently, several FCL methods have leveraged generative models to replace real examples in memory with synthetic data. FedCIL (qi2023FedCIL) employs a GAN with an auxiliary classifier to enable generative replay, thereby mitigating forgetting while aggregating global knowledge across clients. TARGET (zhang2023target) and MFCL (babakniya2024data) introduce data-free knowledge distillation that uses synthetic examples to transfer knowledge from a previously trained global model to client models. LANDER (tran2024lander) builds on this idea by incorporating label text embeddings from pretrained language models as anchors, which enables the generation of more meaningful samples and enhances resistance to forgetting. While effective in mitigating forgetting, these approaches inherit significant drawbacks: training generative models is computationally expensive as image resolution increases and introduces new privacy risks (liu2024generative).
@@
Overall, existing FCL methods face practical challenges in real-world FL deployments due to privacy concerns and resource constraints. Specifically, they often: (1) assume the task identity is known during inference; (2) store exemplars from previous tasks on the server; or (3) train a generative model to synthesize replay samples. In contrast, FedProTIP requires none of these assumptions. It is explicitly designed for task-agnostic inference, where task labels are unavailable, and achieves this without replay buffers, generative models, or auxiliary task classifiers. Instead, FedProTIP leverages lightweight subspace representations for both knowledge retention and task-identity prediction, enabling effective task-agnostic FCL. This focus connects to the broader literature on class-incremental learning (CIL). (kim2022theoretical) provides a theoretical perspective that decomposes the CIL problem into within-task classification and task-identity prediction, showing that both are necessary and sufficient conditions for strong performance. While centralized methods address the challenge of task-agnostic inference via out-of-distribution detection (kim2022theoretical; kim2022continual), per-class classifiers or generative models (zajac2024pec; van2021class), or supervised contrastive learning with nearest-class-mean classifiers (mai2021supervised), none extend naturally to federated environments. Our work is the first to bring task-identity prediction into the FCL paradigm, providing a replay-free, privacy-preserving solution in the challenging task-agnostic CIL setting.
@@ 2.2 Gradient Projection in Continual Learning
Gradient projection methods (zeng2019OWM; farajtabar2020OGD; chaudhry2020orthog-subspace) for continual learning mitigate forgetting by updating model parameters in directions orthogonal to those associated with previous tasks, thereby eliminating the need to store raw data or train generative models. GPM (saha2021GPM) extends these approaches by extracting the bases of low-dimensional subspaces spanned by prior task representations and constraining new gradients to lie orthogonal to these subspaces. A series of follow-up works, including TRGP (lin2022trgp), CUBER (lin2022cuber), SGP (saha2023continual) and DualGPM (liang2023adaptive), relax the orthogonality constraints introduced in GPM to better balance stability and plasticity, which leads to improved continual learning performance. However, directly extending gradient projection methods to federated continual learning (FCL) is nontrivial due to the decentralized nature of data and limited communication budgets. FOT (bakman2024FOT) represents a recent attempt to adapt GPM to the FCL setting by requiring clients to share their raw feature embeddings with a central server to construct gradient subspaces. This approach introduces privacy risks and communication overhead, and its effectiveness is limited to scenarios where task identities are known at inference time – an impractical assumption in many real-world settings. In contrast, FedProTIP achieves state-of-the-art FCL performance using a communication- and computation-efficient projection scheme, without exposing raw embeddings or relying on task labels during inference.
@@ 3.2 Gradient Projection Memory
Gradient projection memory (saha2021GPM) is a replay-free CL scheme that requires storing only a set of core bases $\bm{\Phi}_{l}^{(1:t)}$ extracted from layer-wise activations after fine-tuning the model on $t$ tasks. Specifically, let $\mathbf{W}_{l}^{(t)}\subset\mathbf{W}^{(t)}$ denote the parameters of layer $l$ after training on task $t$, and let $\mathbf{a}_{l}^{(t)}\in\mathbb{R}^{d_{l}\times m}$ represent the input activations to layer $l$ for $m$ training samples $\mathbf{x}^{(t)}$, where $d_{l}$ is the dimensionality of the activations. By applying singular value decomposition (SVD), GPM extracts a set of orthonormal bases $\bm{\Phi}_{l}^{(t)}\in\mathbb{R}^{d_{l}\times r_{l}^{(t)}}$ that span the dominant subspace of task $t$ activations and aggregates them with the existing bases $\bm{\Phi}_{l}^{(1:t-1)}$. During training on the (t+1)(t{+}1)-th task, the parameter update for layer $l$, denoted $\Delta\mathbf{W}_{l}^{(t+1)}$, is projected onto the orthogonal complement of the subspace spanned by $\bm{\Phi}_{l}^{(1:t)}$,
@@ 4 Methodology
While GPM has proven effective in centralized continual learning, extending it to federated settings poses major challenges. FOT (bakman2024FOT) offers an early adaptation by having clients share layer-wise intermediate activations, which the server uses to extract core bases. However, this approach raises significant privacy concerns, as such activations can be exploited in gradient inversion attacks (geiping2020inverting; chen2024recovering). It also introduces substantial communication overhead due to the high dimensionality of the transmitted activations.
@@ 4.2 Extracting Local Core Bases
After completing projected gradient descent-based local training, clients follow the standard federated learning protocol by sending local model updates to the server for aggregation. The server then broadcasts the updated global model $\mathbf{W}^{(t)}$ to all clients, which proceed to perform local core basis extraction. Following the GPM strategy (saha2021GPM), each client $k$ samples $m$ examples from its local dataset $\mathcal{D}_{k}^{(t)}$, feeds them through the received model $\mathbf{W}^{(t)}$, and collects layer-wise intermediate activations. To reduce storage and communication overhead, we introduce a random activation sampling step: from the $m$ activations, a smaller subset of size $m^{s}\!\!\ll\!\!m$ is randomly selected, yielding $\mathbf{a}_{l}^{(t)}\in\mathbb{R}^{d_{l}\times m^{s}}$. These activations are projected onto the orthogonal complement of the previously learned feature subspace by subtracting their component along the existing bases,
@@ 4.4 Task Identity Inference via Subspace Relevance
In continual learning, the feature extractor (encoder) is fine-tuned across sequential tasks, while the decision layer $f_{1:t}(\cdot)$ expands as new tasks are introduced. For example, in class-incremental settings, the output dimensionality of the softmax layer grows with the number of classes. Prior works (saha2021GPM; bakman2024FOT) assume that the task identity $\tau$ is known at inference time, so predictions can be routed through the corresponding decision head $f_{\tau}(\cdot)$. However, this assumption is unrealistic in real-world deployments, where task labels are typically unavailable (kim2022theoretical).
@@ 5 Experiments
We evaluate FedProTIP on three standard continual learning benchmarks: CIFAR100 and ImageNet-R (hendrycks2021many) for class-incremental learning, and DomainNet (peng2019moment) for domain-incremental learning. Comparisons are made against six baselines: FedAvg (fedavg), GLFC (Dong2022GLFC), LGA (dong2023LGA), TARGET (zhang2023target), FOT (bakman2024FOT), and LANDER (tran2024lander). Following (yurochkin2019bayesian), we simulate non-IID client distributions by sampling data partitions via a Dirichlet distribution with varying concentration parameter $\alpha$ (lower $\alpha$ implies greater heterogeneity). All methods use a ResNet-18 backbone pretrained on ImageNet-1K (he2016deep) and fine-tuned on each benchmark dataset. Additional results using ResNets and ViTs trained from scratch are provided in Appendix A.1.
Following prior work (chaudhry2018riemannian), we evaluate performance using two standard metrics: average accuracy (ACC) and forgetting (FT), defined as
@@ 5.2 Robustness under Data Heterogeneity and Forgetting
As shown in Table 1, FedProTIP consistently outperforms all baselines across various values of the Dirichlet concentration parameter $\alpha$, which controls the degree of data heterogeneity (larger $\alpha$ corresponds to more IID-like partitions). As heterogeneity increases (i.e., smaller $\alpha$), client drift (karimireddy2020scaffold) exacerbates catastrophic forgetting, degrading the performance of most methods. Despite this, FedProTIP remains robust. For example, on CIFAR100, while competing baselines like FOT and LANDER experience accuracy drops of 12% and 15%, respectively, when moving from IID to $\alpha=0.2$, FedProTIP sees only a 6% decline. Moreover, it sustains high accuracy while keeping forgetting near zero, demonstrating strong resilience to heterogeneous client updates. Beyond class-incremental learning, we evaluate robustness under overlapping task semantics using DomainNet in a domain-incremental setting, where all tasks share the same label space. In the task-agnostic scenario, both the baselines and FedProTIP (-t) use a single shared classifier. Under this setup, FedProTIP (-t) consistently outperforms all competing methods across varying levels of heterogeneity, demonstrating benefits that extend beyond task identity prediction by enabling more robust representation learning. In contrast, FedProTIP maintains separate classifiers for each task and uses task prediction to route test inputs. Although this approach yields slightly lower accuracy than FedProTIP (-t), which uses a shared classifier, FedProTIP remains competitive – it achieves the lowest forgetting across all settings and delivers accuracy second only to our FedProTIP (-t).
@@ 5.3 Ablation Studies
We investigate whether task-agnostic continual learning methods originally developed for centralized settings can be adapted to federated scenarios. Specifically, we consider the replay-free PEC (zajac2024pec), which assigns a separate classifier to each class, and the replay-based SCR (mai2021supervised), which uses a nearest-class-mean classifier. For SCR, we average class prototypes across clients at inference time while keeping replay data local, and use FedAvg for model aggregation. As shown in Table 5, both methods yield significantly lower accuracy under task-agnostic inference. While SCR outperforms PEC, it relies on clients retaining local data, which may violate privacy constraints. We also evaluate LODE (liang2023loss), which decouples intra- and inter-task losses to implicitly support task-agnostic inference. Applied to the generative replay method TARGET, LODE offers only marginal gains ($2.50\%$), suggesting that naive loss decoupling is insufficient for robust performance in federated settings.
@@ A.1 Results on Different Models
We also evaluate our method on a different backbone, pre-trained ViT-B/16 (steiner2021augreg), as reported in Table 6. In this setting, CIFAR100 images ($3\times 32\times 32$) are resized to $224\times 224$ to match the ViT input resolution. We set the number of local epochs to $5$ and perform $20$ global rounds per task. For Target and Lander, we observed that generating synthetic images at the native $32\times 32$ resolution and subsequently applying resizing augmentation yields better performance, and we adopt this strategy in our experiments.